# Linear colorings of graphs

### v1

@@ 7 Conclusion
In this paper we explored various aspects of the linear chromatic number of graphs. In particular, we gave improved bounds for several graph classes. We recall that most our results in this direction are listed in Table 1. We investigated the subgraph obstructions to bounded linear chromatic number and provided two algorithmic results: a NP-completeness proof and an FPT algorithm. So far, the conjecture that motivated this research is still open. Let us restate it below.
@@ Acknowledgements.
The authors thank Jana Masaříková and Wojciech Nadara for allowing them to include Section 2.2 and its proof. This work was supported in part by the Slovenian Research and Innovation Agency (I0-0035, research program P1-0285 and research projects J1-3003, J1-4008, J1-4084, J1-60012, J5-4596, N1-0209, and N1-0370, and by the research program CogniCom (0013103) at the University of Primorska. J.-F. Raymond was partially supported by the ANR project GRALMECO (ANR-21-CE48-0004).
@@ Appendix A Pseudogrids
The lower bound is Section 3. The upper bound can be obtained by induction on $k$ as follows, with $c^{\prime}=4$. The base case $k=2$ is trivial since the two coloring numbers are equal. When $k>2$, we pick a set of vertices $X$ consisting of the middle row and the middle column of $\boxplus_{k}$ (if $k$ is even, we choose one of the two rows and columns arbitrarily). This set contains $12k-$ 1 vertices, to which we assign a unique color each. Observe that each connected component of $\boxplus_{k}-X$ is a copy of $\boxplus_{k^{\prime}}$ for some $k^{\prime}\leqslant(k-1)/2$. By induction, there is a centered coloring of each such component with at most $c^{\prime}k^{\prime}$ colors. Fix such a coloring, with the same set of colors of each of the 4 components. We claim that this together with the aforementioned coloring of $X$ is a centered coloring of $\boxplus_{k}$ with at most $c^{\prime}k$ colors. The colors bound holds since we use at most $2k$ colors for $X$ and at most $c^{\prime}k/2$ for $\boxplus_{k}-X$, so at most $c^{\prime}k$ in total as $c^{\prime}=4$. It is a centered coloring because any subgraph of $\boxplus_{k}$ either is included in a single component of $G-X$, in which case it has a uniquely colored vertex, by induction, or the subgraph contains a vertex of $X$, which has a unique color by definition.
We now prove the second part of the statement. We may assume that $k\geqslant 3$. By Section 3 we have $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant ck$. Here we will use the following equivalent (see, e.g., [Die25]) definition of a minor: $H^{\prime}$ is a minor of $H$ if there is a function mapping vertices of $H^{\prime}$ to disjoint subsets of $H$, each inducing a connected subgraph, and such that adjacent vertices of $H^{\prime}$ are mapped to subsets that are connected by an edge in $G$. So in the case we study, let $f$ be such a mapping from $V(\boxplus_{k})$ to vertex subsets of $G$.
@@
Suppose first that for every $v\in V(\boxplus_{k})$, the set $f(v)$ contains at most two vertices. Then we can use a coloring of $\boxplus_{k}$ with $c^{\prime}k$ colors as described above and duplicate each color in order to color the vertices in $G$. For instance, if $v\in V(\boxplus_{k})$ receives some color $x$, we assign colors $(0,x)$ and $(1,x)$ to the two vertices of $f(v)$. As noted above, the image of $f$ defines a partition of $V(G)$. This shows that in this case that $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant 2c^{\prime}k$ and we are done.
@@
Then we can get a centered coloring of $G$ as follows: the vertices that also exist in $W$ receive the same color as in this graph and each maximal path of $G-V(W)$ is colored with the same set of at most $\left\lceil\log s\right\rceil$ colors as in Section 2.2. This is indeed a centered coloring, as every subgraph of $G$ either is a subgraph of one of the aforementioned paths of degree 2 vertices, in which case it has a center by the choice of the colors in this path, or it contains a vertex of $W$, and the center exists by the choice of the coloring of $W$. We get $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant 2c^{\prime}k+\left\lceil\log s\right\rceil$. As noted above $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant ck$ and $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant\left\lceil\log s\right\rceil$, so the ratio $\operatorname{\chi_{\mathrm{cen}}}(G)/\operatorname{\chi_{\mathrm{lin}}}(G)$ is bounded from above by a constant. ∎
@@ Appendix B Obstructions for linear chromatic number 3
[…]
@@
[…]

### v2

@@ 7 Conclusion
In this paper we explored various aspects of the linear chromatic number of graphs. In particular, we gave improved bounds for several graph classes. We recall that most of our results in this direction are listed in Table 1. We investigated the subgraph obstructions to bounded linear chromatic number and provided two algorithmic results: an NP-completeness proof and an FPT algorithm. So far, the conjecture that motivated this research is still open. Let us restate it below.
@@ Acknowledgements.
The authors thank Jana Masaříková and Wojciech Nadara for allowing them to include Section 2.2 and its proof. This work was supported in part by the Slovenian Research and Innovation Agency (I0-0035, research program P1-0285 and research projects J1-3003, J1-4008, J1-4084, J1-60012, J1-70035, J1-70046, J5-4596, and N1-0370), and by the research program CogniCom (0013103) at the University of Primorska. J.-F. Raymond was partially supported by the ANR project GRALMECO (ANR-21-CE48-0004).
@@ Appendix A Pseudogrids
The lower bound is Section 3. The upper bound can be obtained by induction on $k$ as follows, with $c^{\prime}=4$. The base case $k=2$ is trivial since the two coloring numbers are equal. When $k>2$, we pick a set of vertices $X$ consisting of the middle row and the middle column of $\boxplus_{k}$ (if $k$ is even, we choose one of the two rows and columns arbitrarily). This set contains 2k12k-1 vertices, to which we assign a unique color each. Observe that each connected component of $\boxplus_{k}-X$ is a copy of $\boxplus_{k^{\prime}}$ for some $k^{\prime}\leqslant(k-1)/2$. By induction, there is a centered coloring of each such component with at most $c^{\prime}k^{\prime}$ colors. Fix such a coloring, with the same set of colors of each of the 4 components. We claim that this, together with the aforementioned coloring of $X$, is a centered coloring of $\boxplus_{k}$ with at most $c^{\prime}k$ colors. The colors bound holds since we use at most $2k$ colors for $X$ and at most $c^{\prime}k/2$ for $\boxplus_{k}-X$, so at most $c^{\prime}k$ in total as $c^{\prime}=4$. It is a centered coloring because any subgraph of $\boxplus_{k}$ either is included in a single component of GXG-X, in which case it has a uniquely colored vertex, by induction, or the subgraph contains a vertex of $X$, which has a unique color by definition.
We now prove the second part of the statement. We may assume that $k\geqslant 3$. By Section 3 we have $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant ck$. Here we will use the following equivalent (see, e.g., [DIE02]) definition of a minor: $H^{\prime}$ is a minor of $H$ if there is a function mapping vertices of $H^{\prime}$ to disjoint subsets of $H$, each inducing a connected subgraph, and such that adjacent vertices of $H^{\prime}$ are mapped to subsets that are connected by an edge in $G$. So in the case we study, let $f$ be such a mapping from $V(\boxplus_{k})$ to vertex subsets of $G$.
@@
Suppose first that for every $v\in V(\boxplus_{k})$, the set $f(v)$ contains at most two vertices. Then we can use a coloring of $\boxplus_{k}$ with $c^{\prime}k$ colors as described above and duplicate each color in order to color the vertices in $G$. For instance, if $v\in V(\boxplus_{k})$ receives some color $x$, we assign colors $(0,x)$ and $(1,x)$ to the two vertices of $f(v)$. As noted above, the image of $f$ defines a partition of $V(G)$. This shows that in this case $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant 2c^{\prime}k$ and we are done.
@@
Then we can get a centered coloring of $G$ as follows: the vertices that also exist in $W$ receive the same color as in this graph and each maximal path of GV(W)G-V(W) is colored with the same set of at most $\left\lceil\log s\right\rceil$ colors as in Section 2.2. This is indeed a centered coloring, as every subgraph of $G$ either is a subgraph of one of the aforementioned paths of degree 2 vertices, in which case it has a center by the choice of the colors in this path, or it contains a vertex of $W$, and the center exists by the choice of the coloring of $W$. We get $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant 2c^{\prime}k+\left\lceil\log s\right\rceil$. As noted above $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant ck$ and $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant\left\lceil\log s\right\rceil$, so the ratio $\operatorname{\chi_{\mathrm{cen}}}(G)/\operatorname{\chi_{\mathrm{lin}}}(G)$ is bounded from above by a constant. ∎
@@ Appendix B Obstructions for linear chromatic number 3
[Uncaptioned image] For the graph $F_{2}$, labeled as in the figure on the left, we can make the following successive assumptions and implications. 1. $\varphi(c)=0$, $\varphi(d)=1$, and $\varphi(e)=2$ because the vertices form a triangle and we can choose the values by symmetry; 2. $\varphi(f)=0$ or $\varphi(g)=0$ because of the path $\mathit{fdeg}$; suppose without loss of generality that $\varphi(f)=0$; 3. $\varphi(b)=2$ because of the path $\mathit{fdcb}$; 4. $\varphi(a)=1$ because of the path $\mathit{abce}$. Then, the path $\mathit{fdecba}$ has no center, a contradiction.
@@
[Uncaptioned image] For the graph $F_{5}$, labeled as in the figure on the left, note that abcdefg is a $P_{7}$, so it has to have a unique color on $d$, as any other choice of a center would leave a path on at least 4 vertices to be colored with 2 colors. We can make the following successive assumptions and implications. 1. $\varphi(d)=0$ (by symmetry of the colors) and color 0 does not appear on the paths $\mathit{abc}$ and $\mathit{efg}$; 2. $\varphi(h)=0$ because of the path $\mathit{hefg}$ and since color 0 does not appear on $\mathit{efg}$; 3. $\varphi(e)=1$ by symmetry of the colors 1 and 2; 4. $\varphi(c)=2$ because of the path $\mathit{hedc}$; 5. $\varphi(b)=1$ and $\varphi(a)=2$ since color 0 does not appear on the path $\mathit{abc}$. Then, the path $\mathit{abcdeh}$ has no center, a contradiction.
[Uncaptioned image] For the graph $F_{7}$, labeled as in the figure on the left, we can make the following successive assumptions and implications. 1. $\varphi(e)=0$, $\varphi(f)=1$, $\varphi(g)=2$ because the vertices form a triangle and we can choose the values by symmetry; 2. $\varphi(c)=1$ by symmetry of the colors 1 and 2; 3. $\varphi(a)=\varphi(d)=2$ because of the paths $\mathit{feca}$ and $\mathit{fecd}$, respectively; 4. $\varphi(b)=0$ because of the path $\mathit{bacd}$. Then, the path $\mathit{bacefg}$ has no center, a contradiction.
[Uncaptioned image] Finally, consider the graph $F_{9}$, labeled as in the figure on the left. The three colors appear on the cycle $\mathit{cdfe}$, one of them appears twice on non-adjacent vertices of that $C_{4}$. By symmetry of the graph and the colors, we can assume $\varphi(c)=\varphi(f)=0$, $\varphi(d)=1$, and $\varphi(e)=2$. Therefore $\varphi(a)=2$ because of the path $\mathit{acdf}$. Then, the path $\mathit{acef}$has no center, a contradiction.